# Smooth polynomials with several prescribed coefficients

### v3

@@ 1. Introduction
In recent years, many spectacular results have been obtained on important problems combining some arithmetic properties of the integers and some conditions on their digits in a given basis, see for example [2, 5, 6, 20, 21, 25].
@@
For a given index-set ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ with ${\mathcal{I}}\leq\delta(g)\cdot n$ and $0\in{\mathcal{I}}$, it follows from [2, 25], that the number of prime numbers with prescribed digits $\alpha_{i}\in\{0,1,\dots,g-1\}$ ($i\in{\mathcal{I}}$) with $\alpha_{0}\neq 0$ satisfies
@@
Ha [12] investigated the same problem over function fields. Namely, let $\mathbb{F}_{q}$ be the finite field of $q$ elements, and consider the polynomial ring $\mathbb{F}_{q}[t]$ over $\mathbb{F}_{q}$. Let ${\mathcal{P}}(n)$ be the set of monic irreducible polynomials of degree $n$. It is well-known, that the number $\pi_{q}(n)=\#{\mathcal{P}}(n)$ of such irreducibly polynomials satisfies
@@
see e.g. [24, Lemma 4]. Improving earlier result on coefficients of irreducible polynomials [13, 24, 27], Ha showed that the number of irreducible polynomials with preassigned coefficients is proportional to $\pi_{q}(n)$. In particular, let ${\mathcal{M}}(n)$ be the set of monic polynomials of degree $n$ and for an index-set ${\mathcal{I}}\subset\{0,1\dots,n-1\}$ and $\alpha_{i}\in\mathbb{F}_{q}$ ($i\in{\mathcal{I}}$) let $\mathcal{J}$} be the set of monic polynomials of degree $n$ with $i$th coefficient $\alpha_{i}$
@@
then for $\delta\leq 1/4$ we have
@@
$\mathfrak{G}=\left\{\begin{array}[]{cl}1&\text{if $0\not\in{\mathcal{I}}$},\\ 1+\frac{1}{q-1}&\text{if $0\in{\mathcal{I}}$ and $\alpha_{0}\neq 0$},\\ 0&\text{if $0\in{\mathcal{I}}$ and $\alpha_{0}=0$}.\end{array}\right$.
In particular, we have the expected frequency of irreducible polynomials with prescribed coefficients, $({\mathcal{P}}(n)\cap{\mathcal{J}})\sim\frac{\mathfrak{G}q^{n-\#{\mathcal{I}}}}{n}$, if $q\rightarrow\infty$ or $\delta=\#{\mathcal{I}}/n\rightarrow 0$.
@@
if $\sqrt{n\log n}=O(m)$. Moreover, the error terms in (3) and (5) are optimal.
@@
(8) $\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+O\left(\exp\left(O_{\varepsilon}\left(\frac{u\log^{2}(u+1)}{m^{2}}+\frac{1}{u}\right)\right)\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\right)\right)\\ &\qquad+O\left(\frac{q^{n}}{q^{\#{\mathcal{I}}}}m^{2}n^{1/2}q^{-\eta n}\right),\end{split}$
@@
Theorem 1 gives bounds on $({\mathcal{S}}(n,m)\cap{\mathcal{J}})$. However, in a shorter range for $m$ instead of (7) we also have an asymptotic result.
@@
(10) $\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\left(1+O\left(\frac{\log(u+1)}{m}\right)\right)\right)\\ &\qquad+O\left(\frac{q^{n}}{q^{\#{\mathcal{I}}}}m^{2}n^{1/2}q^{-\eta n}\right),\end{split}$
@@
$0<\varepsilon<1/4$. Let ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ and $\alpha_{i}\in\mathbb{F}_{q}$ ($i\in{\mathcal{I}}$).
$m\leq n/100,\quad\text{that is,}\quad u\geq 100.$
@@
$\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+O\left(\exp\left(O_{\varepsilon}\left(\frac{u\log^{2}(u+1)}{m^{2}}+\frac{1}{u}\right)\right)\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\right)\right)\\ &\quad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda(n,m,{\mathcal{I}},\kappa)+O\left(nq^{\left(-\frac{1}{2}+\varepsilon\right)n+3\kappa+\frac{3}{2}}e^{\varepsilon n}\right)\right)+O\left(m^{2}n^{1/2}q^{n-\kappa/2}\right),\end{split}$
@@
Again, if $m$ is in the shorter range (9) instead of (13), a stronger result can be obtained for $({\mathcal{S}}(n,m)\cap{\mathcal{J}})$.
@@
$\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\left(1+O\left(\frac{\log(u+1)}{m}\right)\right)\right)\\ &\quad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda(n,m,{\mathcal{I}},\kappa)+O\left(\kappa q^{\left(-\frac{1}{2}+\varepsilon\right)n+3\kappa+\frac{3}{2}}e^{\varepsilon\kappa}\right)\right)+O\left(m^{2}n^{1/2}q^{n-\kappa/2}\right),\end{split}$
@@
Theorems 1 and 4 give asymptotics to the number smooth polynomials with several prescribed coefficients. Recently, Kuttner and Wang [18], using different techniques, have considered the exact number of smooth polynomials of degree $n$ with given trace, that is, when ${\mathcal{I}}=\{n-1\}$.
In a different perspective, Gorodetsky [9] also investigated the number of smooth polynomials with three prescribed coefficients in characteristic 2 and 5 and reveal, after normalization, a cyclic property of their number in terms of $n$.
Hauck and Shparlinski [14] has also considered similar problems over $\mathbb{Z}$ by proving the existence of smooth integers (that is, integers without large prime factor) with few nonzero binary digits.
The proof of Theorem 4 uses the circles method. The major arc analysis is based on character sum estimates on smooth polynomials, see Section 4, and Bourgain’s argument [2], applied for polynomials over finite fields by Ha [12], see Section 6.
@@ 2. Outline of the proof
Let $\mathbb{K}_{\infty}$ be the set of formal power series
@@
where $k$ is the larges index so that $x_{k}\neq 0$.
@@
For a nonempty subset ${\mathcal{I}}\subset\{0,1,\dots n-1\}$ fix $\alpha_{i}\in\mathbb{F}_{q}[t]$ for $i\in{\mathcal{I}}$. Let $\mathcal{J}$} defined by (2) and write
@@
We investigate this integral by the circle method. To do so, let $\ell$ be a positive parameters to be fixed later such that
$\ell+\kappa<n$
@@ 3.2. Arithmetically distributed relations
see [15, Theorem 8.6] and [17, Lemma 1.1]. We also have $G_{\ell}=q^{\ell}$ and
@@
We have the following lower bound on $\Phi(g)$, see [12, Lemma 2.3],
(20) $\Phi(H)\gg\frac{|g|}{\log_{q}(\deg g)+1}\quad\text{for }g\neq 0.$
@@
The following result, see [17, Theorem 1.3], follows from the Riemann hypothesis for the $L$-functions of algebraic curves over finite fields proved by Weil [28].
@@
with $L(s,\chi)=P_{\chi}(q^{-s})$ and $|\alpha_{i}|=\sqrt{q}$ with at most one exceptional $\alpha_{i}$ with $|\alpha_{i}|=1$.
@@
$3.$ Exponential sums
We have, see [16, Theorem 3.7].
@@
For the following result see [16, Theorem 6.1].
@@
We can prove the following results on polynomial Gauss sums.
@@
$=q^{\ell}\Phi(g)\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}a_{1}\equiv a_{2}\bmod{\mathcal{R}}_{g}\\ a_{1}\equiv a_{2}\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}((a-b)/g)=q^{\ell}\Phi(g)^{2}.$
For the following estimate see [22, Lemma 4].
@@ 3.4. Dickman’s $\rho$ function
We need the following result on Dickman’s $\rho$ function which follows directly from [26, Lemma III. 5.11] and [26, Corollary III. 5.15].
@@
Uniformly over $0\leq v\leq u$, we have
@@ 3.5. Further preliminaries
The following lemma of Hayes [16, Theorem 4.3] is an analogue of a well-known result of Dirichlet in the theory of Diophantine approximation.
@@
For any $\varepsilon>0$ there is a constant $c(\varepsilon)>0$ such that
$\tau(f)\leq c(\varepsilon)|f|^{(2+\varepsilon)/\log\deg f}.$
@@
$\frac{\tau(f)}{\exp(\lambda\deg f)}\leq\prod_{\deg g_{i}\leq 1/\lambda}\frac{e_{i}+1}{1+\lambda e_{i}\deg g_{i}}.$
As there are at most $q^{1/\lambda}$ irreducible polynomial of degree at most $1/\lambda$, we get
@@
Choosing
$\lambda=\frac{2\log q}{\log\deg f},$
we get that the logarithm of (21) is
$\log\frac{\tau(f)}{\exp\left(\frac{2\log q\deg f}{\log\deg f}\right)}\leq(\deg f)^{1/2}\log\frac{\log\deg f}{2\log q}\leq\varepsilon\frac{\log q\deg f}{\log\deg f}$
@@ 4. Distribution of smooth polynomials
    $\left|\sum_{f\in S(n,m)}\chi(f)\right|\leq q^{\frac{n}{2}+\frac{\log\log(\ell+1+\deg g)}{\log(\ell+1+\deg g)}}e^{39q\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}}$
@@
For $r\geq R$ we use (23) in (22), while $r<R$ we use the trivial bound (24). Namely, (23) yields that for $r\geq R$, we have that the coefficient of $z^{r}$ in (22) is bounded in absolute value by
@@
(27) $\log T(z,\chi;m)\leq\sum_{1\leq r<R}\frac{2}{r}(q|z|)^{r}+\sum_{r\geq R}\frac{\ell+1+\deg g}{r}(q^{1/2}|z|)^{r}.$
@@
$\log T(z,\chi;m)$ $\leq 13(\ell+1+\deg g)^{2}q^{-(1/2+\delta)(R-1)}+\frac{\ell+1+\deg g}{R}\cdot\frac{q^{-\delta R}}{1-q^{-\delta}}$
@@
$\log T(z,\chi;m)$ $\leq(\ell+1+\deg g)^{(1-2\delta)}\left(39q^{1/2}+\frac{1}{R\delta\log q}\right)$
$\leq(\ell+1+\deg g)^{(1-2\delta)}\left(39q^{1/2}+\frac{1}{\delta\log(\ell+1+\deg g)}\right).$
Let $C_{Z}$ denote the circle centered at 0 with radius $Z$. Then, by the Cauchy’s integral formula, we have
@@
$\leq\exp\left((\ell+1+\deg g)^{(1-2\delta)}\left(39q^{1/2}+\frac{1}{\delta\log(\ell+1+\deg g)}\right)\right)q^{(1/2+\delta)n}.$
@@
and observe that it satisfies (29) by assumption. If $\ell+1+\deg g>10^{4}$, then $\delta\leq 1/4$, and also, $\delta\leq\frac{1}{\log q}$. Then
$(\ell+1+\deg g)^{(1-2\delta)}\left(39q^{1/2}+\frac{1}{\delta\log(\ell+1+\deg g)}\right)$
$=$ $\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}\left(39q^{1/2}+\frac{1}{\log\log(\ell+1+\deg g)}\right)$
@@
$\log T(z,\chi;m)\leq\sum_{r\geq 1}\frac{\ell+1+\deg g}{r}(q^{-\varepsilon})^{r}\leq(\ell+1+\deg g)\cdot\frac{q^{-\varepsilon}}{1-q^{-\varepsilon}}\leq\varepsilon(\ell+1+\deg g)$
@@
$\sum_{f\in S(n,m)}\chi(f)\ll q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon(\ell+\deg g)},$
@@
If $q>(1+1/\varepsilon)^{1/\varepsilon}$, then we apply (ii) of Lemma 17 directly. Otherwise,
@@
$\{f\in{\mathcal{S}}(n,m):f\equiv b\mod{\mathcal{R}}_{\ell}\}=\frac{1}{q^{\ell}}\Psi(n,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon\ell}\right),$
@@
$\frac{1}{q^{\ell}}\sum_{\chi\neq\chi_{0}}\overline{\chi(b)}\sum_{f\in{\mathcal{S}}(n,m)}\chi(f)\ll q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon\ell}.$

### v4

@@ 1. Introduction
In recent years, many spectacular results have been obtained on important problems combining some arithmetic properties of the integers and some conditions on their digits in a given basis, see for example [2, 6, 9, 10, 17, 26, 27, 31, 33].
@@
Then, there is a $\delta(g)\in(0,1)$, that for a given index-set ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ with ${\mathcal{I}}\leq\delta(g)\cdot n$ and $0\in{\mathcal{I}}$, the number of prime numbers with prescribed digits $\alpha_{i}\in\{0,1,\dots,g-1\}$ ($i\in{\mathcal{I}}$) with $\alpha_{0}\neq 0$ satisfies
@@
Ha [18] investigated the same problem over function fields. Namely, let $\mathbb{F}_{q}$ be the finite field of $q$ elements, and consider the polynomial ring $\mathbb{F}_{q}[t]$ over $\mathbb{F}_{q}$. Let ${\mathcal{P}}(n)$ be the set of monic irreducible polynomials of degree $n$. It is well-known, that the number $\pi_{q}(n)=\#{\mathcal{P}}(n)$ of such irreducible polynomials satisfies
@@
see e.g. [30, Lemma 4]. Improving earlier results on coefficients of irreducible polynomials [19, 30, 34], Ha showed that the number of irreducible polynomials with preassigned coefficients is proportional to $\pi_{q}(n)$. In particular, let ${\mathcal{M}}(n)$ be the set of monic polynomials of degree $n$ and for an index-set ${\mathcal{I}}\subset\{0,1\dots,n-1\}$ and $\alpha_{i}\in\mathbb{F}_{q}$ ($i\in{\mathcal{I}}$) let $\mathcal{J}$} be the set of monic polynomials of degree $n$ with $i$th coefficient $\alpha_{i}$
@@
then for $\delta<1/4$ we have
@@
$\mathfrak{G}=\left\{\begin{array}[]{cl}1&\text{if $0\not\in{\mathcal{I}}$},\\ 1+\frac{1}{q-1}&\text{if $0\in{\mathcal{I}}$ and $\alpha_{0}\neq 0$},\\ 0&\text{if $0\in{\mathcal{I}}$ and $\alpha_{0}=0$}\end{array}\right$.
and the implied constants are absolute. In particular, we have the expected frequency of irreducible polynomials with prescribed coefficients, $({\mathcal{P}}(n)\cap{\mathcal{J}})\sim\frac{\mathfrak{G}q^{n-\#{\mathcal{I}}}}{n}$, if $q\rightarrow\infty$ or $\delta=\#{\mathcal{I}}/n\rightarrow 0$.
@@
if
(6) $\sqrt{n\log n}=O(m)$
see [16, Corollary 1.6.]. Moreover, the error terms in (3) and (5) are optimal.
@@
$\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+O\left(\exp\left(O_{\varepsilon}\left(\frac{u\log^{2}(u+1)}{m^{2}}+\frac{1}{u}\right)\right)\frac{\delta^{-1}\log\delta^{-1}}{q^{C/\delta}}\right)\right)\\ &\qquad+O\left(\frac{q^{(1-\eta)n}}{q^{\#{\mathcal{I}}}}\right),\end{split}$
@@
We remark, that the constant $1/24$ is not optimized and arises from technical aspects of the proof; with additional analysis this bound could be improved.
As the first error term in Theorem 1 might be greater than 1, it gives bounds on $({\mathcal{S}}(n,m)\cap{\mathcal{J}})$ in the range (8). However, in a shorter range for $m$ instead of (8), that is, if $\sqrt{n\log n}=O(m)$, we also have an asymptotic result.
@@
(10) $\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\left(1+O\left(\frac{\log(u+1)}{m}\right)\right)\right)\\ &\qquad+O\left(\frac{q^{(1-\eta)n}}{q^{\#{\mathcal{I}}}}\right),\end{split}$
@@
Indeed, under the condition (9), one has $\exp\left(O_{\varepsilon}\left(\frac{u\log^{2}(u+1)}{m^{2}}+\frac{1}{u}\right)\right)=1+O\left(\frac{\log(u+1)}{m}\right)$.
@@
Let $0<\varepsilon<1/4$. Let ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ and $\alpha_{i}\in\mathbb{F}_{q}$ ($i\in{\mathcal{I}}$).
(13) $m\leq n/100,\quad\text{that is,}\quad u\geq 100.$
@@
$\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+O\left(\exp\left(O_{\varepsilon}\left(\frac{u\log^{2}(u+1)}{m^{2}}+\frac{1}{u}\right)\right)\frac{\delta^{-1}\log\delta^{-1}}{q^{C/\delta}}\right)\right)\\ &\quad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda(n,m,{\mathcal{I}},\kappa)+O\left(q^{\left(-\frac{1}{2}+\varepsilon\right)n+3\kappa}\right)\right)+O\left(m^{2}n^{1/2}q^{n-\kappa/2}\right),\end{split}$
@@
Again, if $m$ is in the shorter range (9) instead of (14), an asymptotic result can be obtained for $({\mathcal{S}}(n,m)\cap{\mathcal{J}})$.
@@
$\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}\left(1+\frac{\delta^{-1}\log\delta}{q^{C/\delta}}\left(1+O\left(\frac{\log(u+1)}{m}\right)\right)\right)\\ &\quad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda(n,m,{\mathcal{I}},\kappa)+O\left(q^{\left(-\frac{1}{2}+\varepsilon\right)n+3\kappa}\right)\right)+O\left(m^{2}n^{1/2}q^{n-\kappa/2}\right),\end{split}$
@@
Related works
Theorems 1 and 4 give asymptotics to the number smooth polynomials with several prescribed coefficients.
Cohen [4, 5] investigated the distribution of given factorization type for large field characteristic, i.e. when $q=p^{r}$ with large prime $p$. Using sieving methods one may derive result on smooth polynomials over such fields. Theorems 1 and 4 cover the case of large field size.
Recently, Kuttner and Wang [24], using different techniques, have considered the exact number of smooth polynomials of degree $n$ with given trace, that is, when ${\mathcal{I}}=\{n-1\}$.
Gorodetsky [15] (see also [25]) obtained an asymptotic formula for $\Psi(n,m)$ beyond the range (6), with a different main term that does not involve the Dickman’s $\rho$ function. The proofs of Theorems 1 and 4 rely on a detailed analysis of $\rho$. Incorporating Gorodetsky’s result into our framework would therefore require new ideas, and we leave this direction as an open problem.
Hauck and Shparlinski [20] have also considered similar problems over $\mathbb{Z}$ by proving the existence of smooth integers (that is, integers without large prime factor) with few nonzero binary digits. In addition, Cumberbatch [8] investigated smooth integers with missing digits.
The proof of Theorem 4 uses the circle method. The major arc analysis is based on character sum estimates on smooth polynomials, see Section 4, and Bourgain’s argument [2], applied for polynomials over finite fields by Ha [18], see Section 6.
@@ 2. Outline of the proof
Let $\mathbb{K}_{\infty}$ be the set of formal Laurent series of $1/t$
@@
where $k$ is the largest index so that $x_{k}\neq 0$.
@@
For a nonempty subset ${\mathcal{I}}\subset\{0,1,\dots n-1\}$ fix $\alpha_{i}\in\mathbb{F}_{q}$ for $i\in{\mathcal{I}}$. Let $\mathcal{J}$} defined by (2) and write
@@
We investigate this integral by the circle method. To do so, let $\ell$ be a positive parameter to be fixed later such that
(19) $\ell<\min\{n-\kappa,n/2\}$
@@ 3.2. Arithmetically distributed relations
see [21, Theorem 8.6] and [23, Lemma 1.1]. We also have $G_{\ell}=q^{\ell}$ and
@@
We have the following lower bound on $\Phi(g)$, see [18, Lemma 2.3],
(22) $\Phi(g)\gg\frac{|g|}{\log_{q}(1+\deg g)+1}\quad\text{for }g\neq 0.$
@@
The following result, see [23, Theorem 1.3], follows from the Riemann hypothesis for the $L$-functions of algebraic curves over finite fields proved by Weil [35].
@@
with $L(s,\chi)=P_{\chi}(q^{-s})$ and $|\alpha_{i}|=\sqrt{q}$ or $|\alpha_{i}|=1$.
@@
3.3. Exponential sums
We have, see [22, Theorem 3.7].
@@ 3.3. Exponential sums
For the following result see [22, Theorem 6.1].
@@
We can prove the following results.
@@
$=q^{\ell}\Phi(g)\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}a_{1}\equiv a_{2}\bmod{\mathcal{R}}_{g}\\ a_{1}\equiv a_{2}\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}((a_{1}-a_{2})/g)=q^{\ell}\Phi(g)^{2}.$
For the following estimate see [28, Lemma 4].
@@ 3.4. Dickman’s $\rho$ function
We need the following result on Dickman’s $\rho$ function which follows directly from [32, Lemma III. 5.11] and [32, Corollary III. 5.15].
@@
Uniformly over $0\leq v\leq u$ and $u\geq 3$, we have
@@ 3.5. Further preliminaries
The following lemma of Hayes [22, Theorem 4.3] is an analogue of a well-known result of Dirichlet in the theory of Diophantine approximation.
@@
For any $\varepsilon>0$ we have
$\tau(f)\ll|f|^{(2+\varepsilon)/\log\deg f},$
where the implied constant may depend on $\varepsilon$.
@@
(23) $\frac{\tau(f)}{\exp(\lambda\deg f)}\leq\prod_{\deg g_{i}\leq 1/\lambda}\frac{e_{i}+1}{1+\lambda e_{i}\deg g_{i}}.$
As there are at most $q^{1/\lambda}$ irreducible polynomials of degree at most $1/\lambda$, we get
@@
Put
$\lambda=\frac{2\log q}{\log\deg f}.$
If $\lambda>1$, then the right hand side of (23) is an empty product, thus we get
$\tau(f)\leq\exp(\lambda\deg f)=|f|^{2/\log\deg f}.$
Otherwise the logarithm of (24) is
$\log\frac{\tau(f)}{\exp\left(\frac{2\log q\deg f}{\log\deg f}\right)}\leq(\deg f)^{1/2}\log\frac{\log\deg f}{2\log q}\leq\varepsilon\frac{(\log q)(\deg f)}{\log\deg f}$
@@ 4. Distribution of smooth polynomials
    $\left|\sum_{f\in S(n,m)}\chi(f)\right|\leq q^{\frac{n}{2}+\frac{\log\log(\ell+1+\deg g)}{\log(\ell+1+\deg g)}n}e^{39q\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}}$
@@
For $r\geq R$ we use (26) in (25), while for $r<R$ we use the trivial bound (27). Namely, (26) yields that for $r\geq R$, we have that the coefficient of $z^{r}$ in (25) is bounded in absolute value by
@@
(30) $\left|\log T(z,\chi;m)\right|\leq\sum_{1\leq r<R}\frac{2}{r}(q|z|)^{r}+\sum_{r\geq R}\frac{\ell+1+\deg g}{r}(q^{1/2}|z|)^{r}.$
@@
$\left|\log T(z,\chi;m)\right|$ $\leq 13(\ell+1+\deg g)^{2}q^{-(1/2+\delta)(R-1)}+\frac{\ell+1+\deg g}{R}\cdot\frac{q^{-\delta R}}{1-q^{-\delta}}$
@@
$\left|\log T(z,\chi;m)\right|$ $\leq(\ell+1+\deg g)^{(1-2\delta)}\left(39q^{1/2}+\frac{2}{R\delta\log q}\right)$
$\leq(\ell+1+\deg g)^{(1-2\delta)}\left(39q^{1/2}+\frac{2}{\delta\log(\ell+1+\deg g)}\right).$
Let $C_{Z}$ denote the circle centered at 0 with radius $Z$ traversed anti-clockwise. Then, by the Cauchy’s integral formula, we have
@@
$\leq\exp\left((\ell+1+\deg g)^{(1-2\delta)}\left(39q^{1/2}+\frac{2}{\delta\log(\ell+1+\deg g)}\right)\right)q^{(1/2+\delta)n}.$
@@
and observe that it satisfies (32) by assumption: If $\ell+1+\deg g>10^{4}$, then $\delta<1/4$, and also, $\delta<\frac{1}{\log q}$. Then
$(\ell+1+\deg g)^{(1-2\delta)}\left(39q^{1/2}+\frac{2}{\delta\log(\ell+1+\deg g)}\right)$
$=$ $\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}\left(39q^{1/2}+\frac{2}{\log\log(\ell+1+\deg g)}\right)$
@@
$\left|\log T(z,\chi;m)\right|\leq\sum_{r\geq 1}\frac{\ell+1+\deg g}{r}(q^{-\varepsilon})^{r}\leq(\ell+1+\deg g)\cdot\frac{q^{-\varepsilon}}{1-q^{-\varepsilon}}\leq\varepsilon(\ell+1+\deg g)$
@@
$\sum_{f\in S(n,m)}\chi(f)\ll q^{\left(\frac{1}{2}+\varepsilon\right)n},$
@@
If $q>(1+1/\varepsilon)^{1/\varepsilon}$, then we apply (ii) of Lemma 17 directly by enlarging $\varepsilon$. If $q<(1+1/\varepsilon)^{1/\varepsilon}$ and $\ell+\deg g$ is small in terms of $q$ (and hence in term of $\varepsilon$), then the bound follows. Otherwise,
@@
$\{f\in{\mathcal{S}}(n,m):f\equiv b\mod{\mathcal{R}}_{\ell}\}=\frac{1}{q^{\ell}}\Psi(n,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}\right),$
@@
$\frac{1}{q^{\ell}}\sum_{\chi\neq\chi_{0}}\overline{\chi(b)}\sum_{f\in{\mathcal{S}}(n,m)}\chi(f)\ll q^{\left(\frac{1}{2}+\varepsilon\right)n}.$